OpenStudy (anonymous):

rewrite 7(40+2) as the sum of two products

Parth (parthkohli):

Using distributive property which says \(a(b + c) = (a\times b) + (a \times c)\): \( \color{Black}{\Rightarrow 7(40+2) = (7 \times 40) + (7 \times 2) }\)

OpenStudy (anonymous):

thank you!!!!!!!